2 springs of spring constants K are connectrd in series,what is net spring constant?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear aliina

let the equivaent spring constant is  k

i f the force F is applied at the end of spring then elngation willbe  x1 = F/k1

          and   x2 = F/k2

 

now replace these two spring with single so  elongation in this spring X =F/k

elnogation should be same in both the case  X = X1+x2

                                                              F/k = F/k1  + F/k2

                                                               k=k1k2/(k1+k2)

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin